Circular No. 6918

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bw IN ESO 184-G82 F. Patat, European Southern Observatory (ESO); and A. Piemonte,

Pontificia Universidad Catolica, report on their observations of SN 1998bw with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph; 45-min exposure, range 315–1015 nm, resolution 1.5 nm) on May 23.39 UT at La Silla: “A full reduction of the CCD spectrogram shows no clear traces of He I lines (such as those at 587.6 and 667.8 nm), which were clearly detected in the type-Ib SN 1984L, a few weeks after maximum light (Harkness et al. 1987, Ap.J. 317, 355). Therefore, SN 1998bw should be classified rather as a type-Ic event. Nevertheless, there are some differences between SN 1998bw and known type-Ic supernovae. In particular, the whole spectrum strongly peaks around 540.0 nm, the O I 777.4-nm absorption is much broader than all other absorptions, and the Fe II features are weak, making this object unique. The most prominent features are identified as Ca II H and K, Si II 635.5-nm (possibly blended with C II 658.0-nm), O I 777.4-nm, the Ca II triplet (at 849.8, 854.2, and 866.2 nm), and a number of Fe II lines. The expansion velocities deduced from the minima of Ca II H and K and Si II 635.5-nm are 11 700 and 9100 km/s, respectively.”

SUPERNOVA 1998bu IN NGC 3368 M. Centurion, Instituto de Astrofisica de Canarias; P. Bonifacio, Os-

servatorio Astronomico di Trieste; N. Walton, Isaac Newton Group of Telescopes; and D. King, Royal Greenwich Observatory, report: “A highresolution (∼ 50 000) CCD spectrum of SN 1998bu was obtained by the Extra-solar Planet Observational Research Team with the Utrecht échelle spectrograph on the 4.2-m William Herschel telescope at La Palma on May 18.61 UT. Preliminary reduction reveals two interstellar components of the Ca II doublet (K and H) from M96 at heliocentric radial velocities of +743 km/s (equivalent width of K, EW_{K} = 0.0107 nm), in agreement with the velocity of the stronger component of Na I also reported by Munari et al. on IAUC 6902, and +750 km/s (EW_{K} = 0.0060 nm), which is not observed in Na I. A weak Na I component at +722 km/s is also observed in our spectrum. Five components of Ca II originating in the galactic interstellar medium are observed at heliocentric radial velocities of −26 km/s (EW_{K} = 0.0077 nm), −8 km/s (EW_{K} = 0.0154 nm), +7 km/s (EW_{K} = 0.0016 nm), +48 km/s (EW_{K} = 0.0012 nm), and +53 km/s (EW_{K} = 0.0022 nm). A single component is observed in Na I at −8 km/s, in agreement with Munari et al. (op.cit.).”

1998 May 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green